OpenStudy (anonymous):

Rationalize the denominator of the expression. Then simplify 4/7-sqrt{3}

OpenStudy (goformit100):

 is this the question

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The formula is \[4\div(7-\sqrt{3})\]

OpenStudy (anonymous):

\[\frac{4}{7-\sqrt{3}}\times\frac{7+\sqrt{3}}{7+\sqrt{3}}=\frac{28+4\sqrt{3}}{49-3}\]